what is the difference between resistance and Resistivity

Grade:11

4 Answers

Shyam Chand
10 years ago
Resistance (R) of a conductor is defined as the ratio of the potential difference V across the ends of the conductor to the current I flowing through it. so , R= V/ I. This equation is also called ohm`s law.
Shyam Chand
10 years ago
see basically , Resistance (R) of a conductor is defined as the ratio of the potential difference V across the ends of the conductor to the current I flowing through it. so , R= V/I. This equation is also called ohm`s law. and specific resistance ( rho) or electrical resistivity of the material of a conductor is defined as the resistance offered by unit length and unit area of cross section by a wire of given material of the conductor. so rho = RA/L , WHERE; R= Resistance of that particular conductor A= area of cross section of wire or any conductor L= whole lenght of it.
sumit singh
10 years ago
The electrical resistance of a circuit component or device is defined as the ratio of the voltage applied to the electric current whichflows through it :R = V/I The factor in the resistance which takes into account the nature of the material is the resistivity p=RA/L
mahesh jadi
10 years ago
resistance: 1. the action of opposing something that you disapprove or disagree with he encountered a general feeling of resistance from many citizens; despite opposition from the newspapers he went ahead 2. any mechanical force that tends to retard or oppose motion 3. a material`s opposition to the flow of electric current; measured in ohms 4. the military action of resisting the enemy`s advance the enemy offered little resistance 5. (medicine) the condition in which an organism can resist disease 6. the capacity of an organism to defend itself against harmful environmental agents these trees are widely planted because of their resistance to salt and smog 7. a secret group organized to overthrow a government or occupation force 8. the degree of unresponsiveness of a disease-causing microorganism to antibiotics or other drugs (as in penicillin-resistant bacteria) 9. (psychiatry) an unwillingness to bring repressed feelings into conscious awareness 10. an electrical device that resists the flow of electrical current 11. group action in opposition to those in powe resistivity: a material`s opposition to the flow of electric current; measured in ohms